OpenStudy (warpedkitten):

If the circulation for newspaper 1 is 384 thousand more than the circulation for newspaper​ 2, and their combined circulation is 2512 ​thousand, find the circulation for each newspaper.

OpenStudy (daniel_chernioglo):

alright, so... x + x + 384 = 2512 - 384 -384 2x = 2128 /2 /2 x= 1064 2512-1064=1448 Your answer is: 1064 and 1448

OpenStudy (daniel_chernioglo):

Hopefully this helped :)

OpenStudy (danieldjpon3):

I didn't get it...

OpenStudy (danieldjpon3):

:P

OpenStudy (danieldjpon3):

Let me try and do it.

OpenStudy (danieldjpon3):

btw 2X = 2122

OpenStudy (danieldjpon3):

But the answer is correct, 1061 and 1443

OpenStudy (danieldjpon3):

And x = 1061 not 1064

OpenStudy (danieldjpon3):

So ye @Daniel_Chernioglo

OpenStudy (danieldjpon3):

ye

OpenStudy (danieldjpon3):

...

OpenStudy (daniel_chernioglo):

did you subtract correctly? @danieldjpon3

OpenStudy (danieldjpon3):

Yes

OpenStudy (daniel_chernioglo):

try subtracting 384 from 2512 one more time

OpenStudy (danieldjpon3):

2128

OpenStudy (daniel_chernioglo):

yup now that leaves us with 2x=2128

OpenStudy (danieldjpon3):

I don't get it, this kid told me to loook it up on here

OpenStudy (danieldjpon3):

look on there

OpenStudy (danieldjpon3):

did a super algebra genius make a mistake?

OpenStudy (daniel_chernioglo):

lol, thats practically the same method

OpenStudy (daniel_chernioglo):

but his equation is not the same

OpenStudy (danieldjpon3):

lol ok

OpenStudy (danieldjpon3):

Oh well ok

OpenStudy (daniel_chernioglo):

yup xD

OpenStudy (danieldjpon3):

So there's to ways to get the answer right?

OpenStudy (danieldjpon3):

two*

OpenStudy (daniel_chernioglo):

Im not sure, there might be. But this is the only way ik how to solve this equation

OpenStudy (danieldjpon3):

lol ok

OpenStudy (danieldjpon3):

yes?

OpenStudy (danieldjpon3):

;-;

OpenStudy (danieldjpon3):

If you give me one :D

OpenStudy (danieldjpon3):

XD